The isoelectric point (pI) of a peptide is the pH at which the peptide does not migrate in an electric field. Since the peptide is zwitterionic, there are the same number of positive charges as negative charges on the peptide population. The pI can be estimated fairly accurately (within 0.1 or 0.2 pH units) from the pK values of all the proton dissociable groups in the peptide. Using pK values from the table at the right, estimate the pI value of the following hexapeptide:The isoelectric point (pl) of a peptide is the pH at which the peptide does not migrate in an electric field. Since the pepti
Answer #1

Antwer Phe -Ls- Atp- Cs-Thr -Tr PE-3.7 OH Ph H N OH H Pro- 3-4 NH (CH)4 SH Pra- 8.0 NH2 P=8-2 OH Ple=10. PKa-lo- Sucture will